OpenStudy (steve816):

Question about whether this function can be differentiated using the power rule.

OpenStudy (steve816):

\[\frac{ 1 }{ \sqrt{5x} }\]

OpenStudy (steve816):

I know that this function can be simplified to \[(5x)^{-\frac{ 1 }{ 2 }}\]

OpenStudy (steve816):

But now, I'm stuck

OpenStudy (wckeller):

Then using the power rule, move the power to the front and the reduce the original power by 1

OpenStudy (steve816):

I know, but that doesn't work because the 5 is not a coefficient

OpenStudy (wckeller):

It would look like this: \[-\frac{ 1 }{ 2 } (5x)^\frac{ -3 }{ 2 }\]

OpenStudy (steve816):

Ohh yeah, that is what I thought. So how would I simplify that?

OpenStudy (wckeller):

then this: \[\frac{ -1 }{ 2*\sqrt{5x}^3 }\]

OpenStudy (steve816):

is the cubed inside of the square root?

OpenStudy (wckeller):

Yeah that was my mistake, sorry.

OpenStudy (steve816):

Yaay thank you for the help :D You are actually really smart.

OpenStudy (wckeller):

Lol, thanks.